pls draw a gragh of----


a) kelvin and celcius


b)celsius and fahrenheit


c)fahrenheit and kelvin


I LL SURELY APPROVE

Grade:9

1 Answers

Ronak Jain
13 years ago

Use Relations

K = °C + 273.15

K = (°F + 459.67) ÷ 1.8

°F = (K × 9/5) − 459.67

°C = K − 273.15

Compare With : (Y=mX+C)